# GPT-4 Technical Report

### v5

@@ A.5 Scoring
We synthesized multiple-choice section scores and free-response section scores into overall scores using the best available approximations of the real methodologies: for the SAT, we converted multiple-choice scores into scaled scores using the score calculation chart from an official sample SAT as republished on an SAT prep site Seigel [2020]; for the GRE, we converted multiple-choice scores to the 130-170 scale using the official formula of multiplying accuracy by 40 and adding 130; for the AP exams, we used the score calculators found on a public study site, which are based on the point values from the official AP scoring guidelines from 2019-2020 alb . Percentiles are based on the most recently available score distributions for test-takers of each exam type.
For percentile results on the AMC 10 and 12, since 2022 score distributions are as yet unpublished, we used two official published score distributions from November 2021 for exams A and B, and took the minimum lower percentile of the two and the maximum upper percentile of the two to report an estimated percentile range of America [2023]. Other percentiles were based on official score distributions Edwards [2022] Board [2022a] Board [2022b] for Excellence in Education [2022] Swimmer [2021].
@@ A.8 Example few-shot prompts
In the example prompt below, the task prompt would be replaced by a prompt like an official sample GRE essay task, and the essay response with an example of a high-scoring essay ETS [2022].
@@ Appendix B Impact of RLHF on capability
Exam
Base model
RLHF model
LSAT (MCQ)
67.0 %
72.0 %
SAT EBRW - Reading Portion
92.3 %
90.4 %
SAT EBRW - Writing Portion
90.9 %
84.1 %
SAT Math (MCQ)
91.4 %
86.2 %
Graduate Record Examination (GRE) Quantitative
57.5 %
67.5 %
Graduate Record Examination (GRE) Verbal
87.5 %
90.0 %
USNCO Local Section Exam 2022
51.7 %
63.3 %
AP Art History (MCQ)
72.5 %
66.2 %
AP Biology (MCQ)
98.3 %
96.7 %
AP Calculus BC (MCQ)
66.7 %
57.8 %
AP Chemistry (MCQ)
58.3 %
71.7 %
AP English Language and Composition (MCQ)
55.6 %
51.1 %
AP English Literature and Composition (MCQ)
63.6 %
69.1 %
AP Environmental Science (MCQ)
72.5 %
67.5 %
AP Macroeconomics (MCQ)
83.3 %
76.7 %
AP Microeconomics (MCQ)
90.0 %
76.7 %
AP Physics 2 (MCQ)
62.2 %
71.1 %
AP Psychology (MCQ)
98.0 %
96.0 %
AP Statistics (MCQ)
60.0 %
62.5 %
AP US Government (MCQ)
85.5 %
83.6 %
AP US History (MCQ)
89.1 %
87.3 %
AP World History (MCQ)
94.5 %
98.2 %
MKSAP Questions (MCQ)
77.9 %
74.7 %
AMC 10
28.0 %
24.0 %
AMC 12
20.0 %
32.0 %
Introductory Sommelier (theory knowledge)
90.5 %
92.2 %
Certified Sommelier (theory knowledge)
83.2 %
86.2 %
Advanced Sommelier (theory knowledge)
74.8 %
77.1 %
Average
73.7 %
74.0 %
@@ Appendix C Contamination on professional and academic exams
Exam
Contam
GPT-4 (no vision)
Non-contaminated GPT-4 (no vision)
GPT-4
Non-contaminated GPT-4
Uniform Bar Exam   (MBE+MEE+MPT)
0 %
298 / 400 (~90th)
298 / 400 (~90th)
298 / 400 (~90th)
298 / 400 (~90th)
LSAT
39 %
161 (~83rd)
167 (~95th)
163 (~88th)
169 (~97th)
SAT Evidence-Based Reading & Writing
12 %
710 / 800 (~93rd)
710 / 800 (~93rd)
710 / 800 (~93rd)
710 / 800 (~93rd)
SAT Math
7 %
700 / 800 (~89th)
690 / 800 (~89th)
710 / 800 (~91st)
700 / 800 (~89th)
GRE Quantitative
35 %
157 / 170 (~62nd)
161 / 170 (~75th)
163 / 170 (~80th)
165 / 170 (~85th)
GRE Verbal
25 %
166 / 170 (~97th)
165 / 170 (~96th)
169 / 170 (~99th)
169 / 170 (~99th)
GRE Writing
100 %
4 / 6 (~54th)
N/A
4 / 6 (~54th)
N/A
USABO Semifinal Exam 2020
3 %
87 / 150 (99th - 100th)
87 / 150 (99th - 100th)
87 / 150 (99th - 100th)
87 / 150 (99th - 100th)
USNCO Local Section Exam 2022
5 %
38 / 60
38 / 60
36 / 60
36 / 60
Medical Knowledge   Self-Assessment Program
19 %
75 %
75 %
75 %
75 %
Codeforces Rating
0 %
392 (below 5th)
392 (below 5th)
392 (below 5th)
392 (below 5th)
AP Art History
17 %
5 (86th - 100th)
5 (86th - 100th)
5 (86th - 100th)
5 (86th - 100th)
AP Biology
1 %
5 (85th - 100th)
5 (85th - 100th)
5 (85th - 100th)
5 (85th - 100th)
AP Calculus BC
3 %
4 (43rd - 59th)
4 (43rd - 59th)
4 (43rd - 59th)
4 (43rd - 59th)
AP Chemistry
16 %
4 (71st - 88th)
4 (71st - 88th)
4 (71st - 88th)
4 (71st - 88th)
AP Eng. Lang. and Comp.
79 %
2 (14th - 44th)
N/A
2 (14th - 44th)
N/A
AP Eng. Lit. and Comp.
92 %
2 (8th - 22nd)
N/A
2 (8th - 22nd)
N/A
AP Environmental Science
4 %
5 (91st - 100th)
5 (91st - 100th)
5 (91st - 100th)
5 (91st - 100th)
AP Macroeconomics
9 %
5 (84th - 100th)
5 (84th - 100th)
5 (84th - 100th)
5 (84th - 100th)
AP Microeconomics
2 %
4 (60th - 82nd)
5 (82nd - 100th)
5 (82nd - 100th)
5 (82nd - 100th)
AP Physics 2
12 %
4 (66th - 84th)
4 (66th - 84th)
4 (66th - 84th)
4 (66th - 84th)
AP Psychology
11 %
5 (83rd - 100th)
5 (83rd - 100th)
5 (83rd - 100th)
5 (83rd - 100th)
AP Statistics
13 %
5 (85th - 100th)
5 (85th - 100th)
5 (85th - 100th)
5 (85th - 100th)
AP US Government
24 %
5 (88th - 100th)
5 (88th - 100th)
5 (88th - 100th)
5 (88th - 100th)
AP US History
73 %
4 (74th - 89th)
4 (74th - 89th)
5 (89th - 100th)
5 (89th - 100th)
AP World History
47 %
5 (87th - 100th)
4 (65th - 87th)
4 (65th - 87th)
4 (65th - 87th)
AMC 10
4 %
36 / 150 (10th - 19th)
38 / 150 (14th - 21st)
30 / 150 (6th - 12th)
31 / 150 (7th - 12th)
AMC 12
4 %
48 / 150 (19th - 40th)
50 / 150 (26th - 44th)
60 / 150 (45th - 66th)
62 / 150 (52nd - 68th)
Introductory Sommelier (theory knowledge)
5 %
92 %
92 %
92 %
92 %
Certified Sommelier (theory knowledge)
9 %
86 %
86 %
86 %
86 %
Advanced Sommelier (theory knowledge)
4 %
77 %
77 %
77 %
77 %
Leetcode (easy)
0 %
31 / 41
31 / 41
31 / 41
31 / 41
Leetcode (medium)
0 %
21 / 80
21 / 80
21 / 80
21 / 80
Leetcode (hard)
0 %
3 / 45
3 / 45
3 / 45
3 / 45
@@ Appendix D Contamination on academic benchmarks
Benchmark
GPT-4
GPT-3.5
Contamination
GPT-4 (non-contaminated)
Degradation
MMLU
86.4%
70.0%
~0.6%
GSM-8K
92.0%
57.1%
~1%
HellaSwag
95.3%
85.5%
AI2
96.3%
85.2%
~3.4%
WinoGrande
87.5%
81.6%
~0.9%
HumanEval
67.0%
48.1%
25%
65.58%
-2.12%
DROP (F1)
80.9
64.1
~21%
82.8* (subsample)
0
@@ Appendix F Multilingual MMLU
We translated all questions and answers from MMLU [Hendrycks et al., 2020] using Azure Translate. We used an external model to perform the translation, instead of relying on GPT-4 itself, in case the model had unrepresentative performance for its own translations. We selected a range of languages that cover different geographic regions and scripts, we show an example question taken from the astronomy category translated into Marathi, Latvian and Welsh in Table 13. The translations are not perfect, in some cases losing subtle information which may hurt performance. Furthermore some translations preserve proper nouns in English, as per translation conventions, which may aid performance.
We incorporated the same MMLU prompt as [Rae et al., 2021], the model is instructed that it is an intelligent agent, supplied with the questions and a list of four answer options labelled ‘A-D’, followed by ‘Answer:’. We translate the model instruction, question and answers, however preserve the ‘Answer’ token along with the ‘A-D’ options in English. An example prompt is shown in Table 12. The prompts are composed three-shot, with the three examples picked from the development set. We use three-shot evaluation over the regular five-shot because some languages map to much longer token sequences. Finally we classify the correct answer by picking the A-D token continuation with the highest probability from the model.
@@
[…]
@@ Appendix G Examples of GPT-4 Visual Input
User
What is the sum of average daily meat consumption for Georgia and Western Asia? Provide a step-by-step reasoning before providing your answer.
GPT-4
To find the sum of average daily meat consumption for Georgia and Western Asia, we will follow these steps:
@@
User
Answer question I.1.a. Think step-by-step.
GPT-4
To answer question I.1.a, we need to express the temperature $T(x)$ in every point of the conductive bar, which is characterized by the abscissa x. We are given that $T(0)=T0$ and $T(L)=Tb$.
@@
$T(x)=(Tb-T0)*(x/L)+T0$
This represents a linear temperature profile along the conductive bar, with a slope of $(Tb-T0)/L$.
@@
User
What is unusual about this image?
@@
GPT-4
The unusual thing about this image is that a man is ironing clothes on an ironing board attached to the roof of a moving taxi.
@@
User
Below is part of the InstuctGPT paper. Could you read and summarize it to me?
@@
GPT-4
The InstructGPT paper focuses on training large language models to follow instructions with human feedback. The authors note that making language models larger doesn’t inherently make them better at following a user’s intent. Large models can generate outputs that are untruthful, toxic, or simply unhelpful.
@@
User
Could you explain the process described in Figure 2?
GPT-4
Figure 2 in the InstructGPT paper illustrates the three steps of their method:
@@
User
Can you explain this meme?
GPT-4
This meme is a joke that combines two unrelated things: pictures of the earth from space and chicken nuggets.
@@
User
Can you explain why this is funny. Think about it step-by-step.
GPT-4
The comic is satirizing the difference in approaches to improving model performance between statistical learning and neural networks.
@@ Appendix H System Card
The System Card Mitchell et al. [2019], Green et al. [2022] for GPT-4 is appended to this document.

### v6

@@ A.5 Scoring
We synthesized multiple-choice section scores and free-response section scores into overall scores using the best available approximations of the real methodologies: for the SAT, we converted multiple-choice scores into scaled scores using the score calculation chart from an official sample SAT as republished on an SAT prep site [74]; for the GRE, we converted multiple-choice scores to the 130-170 scale using the official formula of multiplying accuracy by 40 and adding 130; for the AP exams, we used the score calculators found on a public study site, which are based on the point values from the official AP scoring guidelines from 2019-2020 [75]. Percentiles are based on the most recently available score distributions for test-takers of each exam type.
For percentile results on the AMC 10 and 12, since 2022 score distributions are as yet unpublished, we used two official published score distributions from November 2021 for exams A and B, and took the minimum lower percentile of the two and the maximum upper percentile of the two to report an estimated percentile range [76]. Other percentiles were based on official score distributions [77] [78] [79] [80] [81].
@@ A.8 Example few-shot prompts
In the example prompt below, the task prompt would be replaced by a prompt like an official sample GRE essay task, and the essay response with an example of a high-scoring essay [83].
@@ Appendix B Impact of RLHF on capability
Exam Base model RLHF model
LSAT (MCQ) 67.0 % 72.0 %
SAT EBRW - Reading Portion 92.3 % 90.4 %
SAT EBRW - Writing Portion 90.9 % 84.1 %
SAT Math (MCQ) 91.4 % 86.2 %
Graduate Record Examination (GRE) Quantitative 57.5 % 67.5 %
Graduate Record Examination (GRE) Verbal 87.5 % 90.0 %
USNCO Local Section Exam 2022 51.7 % 63.3 %
AP Art History (MCQ) 72.5 % 66.2 %
AP Biology (MCQ) 98.3 % 96.7 %
AP Calculus BC (MCQ) 66.7 % 57.8 %
AP Chemistry (MCQ) 58.3 % 71.7 %
AP English Language and Composition (MCQ) 55.6 % 51.1 %
AP English Literature and Composition (MCQ) 63.6 % 69.1 %
AP Environmental Science (MCQ) 72.5 % 67.5 %
AP Macroeconomics (MCQ) 83.3 % 76.7 %
AP Microeconomics (MCQ) 90.0 % 76.7 %
AP Physics 2 (MCQ) 62.2 % 71.1 %
AP Psychology (MCQ) 98.0 % 96.0 %
AP Statistics (MCQ) 60.0 % 62.5 %
AP US Government (MCQ) 85.5 % 83.6 %
AP US History (MCQ) 89.1 % 87.3 %
AP World History (MCQ) 94.5 % 98.2 %
MKSAP Questions (MCQ) 77.9 % 74.7 %
AMC 10 28.0 % 24.0 %
AMC 12 20.0 % 32.0 %
Introductory Sommelier (theory knowledge) 90.5 % 92.2 %
Certified Sommelier (theory knowledge) 83.2 % 86.2 %
Advanced Sommelier (theory knowledge) 74.8 % 77.1 %
Average 73.7 % 74.0 %
@@ Appendix C Contamination on professional and academic exams
Exam Contam GPT-4 (no vision) Non-contaminated GPT-4 (no vision) GPT-4 Non-contaminated GPT-4
Uniform Bar Exam (MBE+MEE+MPT) 0 % 298 / 400 (~90th) 298 / 400 (~90th) 298 / 400 (~90th) 298 / 400 (~90th)
LSAT 39 % 161 (~83rd) 167 (~95th) 163 (~88th) 169 (~97th)
SAT Evidence-Based Reading & Writing 12 % 710 / 800 (~93rd) 710 / 800 (~93rd) 710 / 800 (~93rd) 710 / 800 (~93rd)
SAT Math 7 % 700 / 800 (~89th) 690 / 800 (~89th) 710 / 800 (~91st) 700 / 800 (~89th)
GRE Quantitative 35 % 157 / 170 (~62nd) 161 / 170 (~75th) 163 / 170 (~80th) 165 / 170 (~85th)
GRE Verbal 25 % 166 / 170 (~97th) 165 / 170 (~96th) 169 / 170 (~99th) 169 / 170 (~99th)
GRE Writing 100 % 4 / 6 (~54th) N/A 4 / 6 (~54th) N/A
USABO Semifinal Exam 2020 3 % 87 / 150 (99th - 100th) 87 / 150 (99th - 100th) 87 / 150 (99th - 100th) 87 / 150 (99th - 100th)
USNCO Local Section Exam 2022 5 % 38 / 60 38 / 60 36 / 60 36 / 60
Medical Knowledge Self-Assessment Program 19 % 75 % 75 % 75 % 75 %
Codeforces Rating 0 % 392 (below 5th) 392 (below 5th) 392 (below 5th) 392 (below 5th)
AP Art History 17 % 5 (86th - 100th) 5 (86th - 100th) 5 (86th - 100th) 5 (86th - 100th)
AP Biology 1 % 5 (85th - 100th) 5 (85th - 100th) 5 (85th - 100th) 5 (85th - 100th)
AP Calculus BC 3 % 4 (43rd - 59th) 4 (43rd - 59th) 4 (43rd - 59th) 4 (43rd - 59th)
AP Chemistry 16 % 4 (71st - 88th) 4 (71st - 88th) 4 (71st - 88th) 4 (71st - 88th)
AP Eng. Lang. and Comp. 79 % 2 (14th - 44th) N/A 2 (14th - 44th) N/A
AP Eng. Lit. and Comp. 92 % 2 (8th - 22nd) N/A 2 (8th - 22nd) N/A
AP Environmental Science 4 % 5 (91st - 100th) 5 (91st - 100th) 5 (91st - 100th) 5 (91st - 100th)
AP Macroeconomics 9 % 5 (84th - 100th) 5 (84th - 100th) 5 (84th - 100th) 5 (84th - 100th)
AP Microeconomics 2 % 4 (60th - 82nd) 5 (82nd - 100th) 5 (82nd - 100th) 5 (82nd - 100th)
AP Physics 2 12 % 4 (66th - 84th) 4 (66th - 84th) 4 (66th - 84th) 4 (66th - 84th)
AP Psychology 11 % 5 (83rd - 100th) 5 (83rd - 100th) 5 (83rd - 100th) 5 (83rd - 100th)
AP Statistics 13 % 5 (85th - 100th) 5 (85th - 100th) 5 (85th - 100th) 5 (85th - 100th)
AP US Government 24 % 5 (88th - 100th) 5 (88th - 100th) 5 (88th - 100th) 5 (88th - 100th)
AP US History 73 % 4 (74th - 89th) 4 (74th - 89th) 5 (89th - 100th) 5 (89th - 100th)
AP World History 47 % 5 (87th - 100th) 4 (65th - 87th) 4 (65th - 87th) 4 (65th - 87th)
AMC 10 4 % 36 / 150 (10th - 19th) 38 / 150 (14th - 21st) 30 / 150 (6th - 12th) 31 / 150 (7th - 12th)
AMC 12 4 % 48 / 150 (19th - 40th) 50 / 150 (26th - 44th) 60 / 150 (45th - 66th) 62 / 150 (52nd - 68th)
Introductory Sommelier (theory knowledge) 5 % 92 % 92 % 92 % 92 %
Certified Sommelier (theory knowledge) 9 % 86 % 86 % 86 % 86 %
Advanced Sommelier (theory knowledge) 4 % 77 % 77 % 77 % 77 %
Leetcode (easy) 0 % 31 / 41 31 / 41 31 / 41 31 / 41
Leetcode (medium) 0 % 21 / 80 21 / 80 21 / 80 21 / 80
Leetcode (hard) 0 % 3 / 45 3 / 45 3 / 45 3 / 45
@@ Appendix D Contamination on academic benchmarks
Benchmark GPT-4 GPT-3.5 Contamination GPT-4 (non-contaminated) Degradation
MMLU 86.4% 70.0% ~0.6% - -
GSM-8K 92.0% 57.1% ~1% - -
HellaSwag 95.3% 85.5% -* - -
AI2 96.3% 85.2% ~3.4% - -
WinoGrande 87.5% 81.6% ~0.9% - -
HumanEval 67.0% 48.1% 25% 65.58% -2.12%
DROP (F1) 80.9 64.1 ~21% 82.8* (subsample) 0
@@ Appendix F Multilingual MMLU
We translated all questions and answers from MMLU [49] using Azure Translate. We used an external model to perform the translation, instead of relying on GPT-4 itself, in case the model had unrepresentative performance for its own translations. We selected a range of languages that cover different geographic regions and scripts, we show an example question taken from the astronomy category translated into Marathi, Latvian and Welsh in Table 13. The translations are not perfect, in some cases losing subtle information which may hurt performance. Furthermore some translations preserve proper nouns in English, as per translation conventions, which may aid performance.
We incorporated the same MMLU prompt as [4], the model is instructed that it is an intelligent agent, supplied with the questions and a list of four answer options labelled ‘A-D’, followed by ‘Answer:’. We translate the model instruction, question and answers, however preserve the ‘Answer’ token along with the ‘A-D’ options in English. An example prompt is shown in Table 12. The prompts are composed three-shot, with the three examples picked from the development set. We use three-shot evaluation over the regular five-shot because some languages map to much longer token sequences. Finally we classify the correct answer by picking the A-D token continuation with the highest probability from the model.
@@
aAkAf EnfflΛ kA aAhΛ ?
A) kArZ pΘ LvFQyA vAtAvrZAcF rcnA krZAyA rΛZΔ cA rg EnfflA asto
B) kArZ aAkAfAtΔ n pΘ LvFQyA mhAsAgrAcA rg EtEbEbt hoto
C) kArZ vAtAvrZ AmΓ HyAnΛ lhAn trglAbF EvKΓ rtΛ
D) kArZ pΘ LvFcΛ vAtAvrZ itr sv rgAnA ADAΣyAnΛ foqΔ n GΛtΛ
@@ Appendix G Examples of GPT-4 Visual Input
User What is the sum of average daily meat consumption for Georgia and Western Asia? Provide a step-by-step reasoning before providing your answer.
GPT-4 To find the sum of average daily meat consumption for Georgia and Western Asia, we will follow these steps:
@@
User Answer question I.1.a. Think step-by-step.
GPT-4 To answer question I.1.a, we need to express the temperature $T(x)$ in every point of the conductive bar, which is characterized by the abscissa x. We are given that $T(0)=T0$ and $T(L)=Tb$.
@@
T(x)=(TbT0)(x/L)+T0T(x)=(Tb-T0)*(x/L)+T0
This represents a linear temperature profile along the conductive bar, with a slope of (TbT0)/L(Tb-T0)/L.
@@
User What is unusual about this image?
@@
GPT-4 The unusual thing about this image is that a man is ironing clothes on an ironing board attached to the roof of a moving taxi.
@@
User Below is part of the InstuctGPT paper. Could you read and summarize it to me?
@@
GPT-4 The InstructGPT paper focuses on training large language models to follow instructions with human feedback. The authors note that making language models larger doesn’t inherently make them better at following a user’s intent. Large models can generate outputs that are untruthful, toxic, or simply unhelpful.
@@
User Could you explain the process described in Figure 2?
GPT-4 Figure 2 in the InstructGPT paper illustrates the three steps of their method:
@@
User Can you explain this meme?
GPT-4 This meme is a joke that combines two unrelated things: pictures of the earth from space and chicken nuggets.
@@
User Can you explain why this is funny. Think about it step-by-step.
GPT-4 The comic is satirizing the difference in approaches to improving model performance between statistical learning and neural networks.
@@ Appendix H System Card
The System Card [84, 85] for GPT-4 is appended to this document.